OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) = Ie^(rt). Juan was given 600 milligrams of medicine which leaves his bloodstream at a rate of 25%. How much of the medicine remains in his system after 6 hours?

OpenStudy (tkhunny):

"25%" is not a rate. 25% per hour? \(A(t) = 600e^{rt}\) We know that \(A(1) = 0.25*600 = 150\) -- This can be used to find the proper value for 'r'. Go!

OpenStudy (anonymous):

I think it should be \[A(6)=0.25\times600=150\]

OpenStudy (anonymous):

and that approximately 150 mg of the medicine would be in his blood after 6 hours. Am I right?

OpenStudy (tkhunny):

Well, that would be part of the problem with your language. You did not address my question about 25%. I suspect the rate is 25% / hour, making A(1) = 150. This equation will allow you to find 'r'. This value for 'r' will allow you to find A(6).

OpenStudy (tkhunny):

Whoops! The exit rate is 25% / hour, not the remainder. This makes A(1) = 600(1 - 25%) = 450. Sorry about that.

OpenStudy (anonymous):

How do i use A(1)=450 to find r?

OpenStudy (tkhunny):

\(A(t) = 600\cdot e^{r\cdot t}\) \(A(1) = 600\cdot e^{r\cdot (1)} = 450\) - -Solve for r.